### v1

# CasTex: Cascaded Text-to-Texture Synthesis via Explicit Texture Maps and Physically-Based Shading

Abstract
This work investigates text-to-texture synthesis using diffusion models to generate physically-based texture maps. We aim to achieve realistic model appearances under varying lighting conditions. A prominent solution for the task is score distillation sampling. It allows recovering a complex texture using gradient guidance given a differentiable rasterization and shading pipeline. However, in practice, the aforementioned solution in conjunction with the widespread latent diffusion models produces severe visual artifacts and requires additional regularization such as implicit texture parameterization. As a more direct alternative, we propose an approach using cascaded diffusion models for texture synthesis (CasTex). In our setup, score distillation sampling yields high-quality textures out-of-the box. In particular, we were able to omit implicit texture parameterization in favor of an explicit parameterization to improve the procedure. In the experiments, we show that our approach significantly outperforms state-of-the-art optimization-based solutions on public texture synthesis benchmarks.
Project page: https://thecrazymage.github.io/CasTex/.
1 Introduction
The creation of 3D assets is a crucial part of modern computer graphics applications, with demand being supported by such industries as movie production, industrial design, and computer gaming, to name a few. High-quality assets should simultaneously meet aesthetic requirements and technical production requirements (i.e., high-quality mesh topology and limited polygon count). As a result, designing a 3D model is a labor-intensive creative task that involves multiple skills and is hard to automate due to the lack of high-quality data and occasionally difficulty in formalizing target criteria.
At the same time, the rapid development of generative models in the visual domain has already tremendously impacted various creative domains and even extends to 3D graphics. The successes in 3D reconstruction [25, 1, 17] and foundation text-to-image model [33, 32] led fully-automated end-to-end 3d asset creation prototypes [29, 23, 20]. Such solutions use the prior knowledge acquired by the diffusion model to generate 3D assets meeting aesthetic requirements. However, the resulting assets come in a volumetric representation and are often incompatible with modern graphics engines. Even with conversion tools for extracting meshes from volumetric representation, the assets end up having overly detailed geometry with occasional high-level geometric inconsistencies (e.g., multiple faces, as in the Janus effect). The aforementioned limitations may be critical when the technical requirements come into play.
Texture synthesis is a more modest attempt to automate 3D asset creation, aiming to create a realistic texture for a given 3D shape. The task assumes that all of the technical requirements regarding the model geometry are already addressed by an artist. To meet modern industry standards, in our work, we specifically consider the synthesis of physically based textures [4]. Such textures present a versatile instrument able to capture a diversity of materials and create realistically looking assets within varying lighting scenarios.
Some early texture synthesis approaches involved the back-projection of synthetic images on 3D shapes. However, such approaches are incompatible with PBR textures as they only allow the capture of a fraction of material properties through RGB colors. As an alternative, we consider an optimization based approach that allows capturing the material properties with back-propagation through the rendering engine. In particular, we continue the line of work studying various applications of Score Distillation Sampling [29].
We aimed to design a zero-shot texture-synthesis solution without relying on auxiliary 3D data and foundation model fine-tuning. However, we note that the texture synthesis setup unveils the limitations of Score Distillation Sampling applied in conjunction with modern diffusion models. Our observations go in line with the results of [41]. We continue their analysis and attribute the effect to the specifics of latent diffusion models.
As an alternative, we consider cascaded diffusion models that guide the images in RGB space rather than latent space. Our solution follows the pipeline of cascaded models. First, we generate a coarse texture using SDS guidance with a low-resolution model. Then we refine the texture using SDS guidance with a super-resolution model. Below, we summarize our main contributions.
  • We study the limitations of Score Distillation Sampling when applied to latent diffusion models. Our analysis indicates that the optimal solutions of the algorithm include images outside of the natural image domain.
  • Motivated by our analysis, we introduce a texture synthesis approach using Score Distillation Sampling with cascaded diffusion models. The proposed approach mimicks the diffusion model architecture and acts in two stages. The first stage generates a coarse texture, while the second stage improves overall quality of the texture obtained at the first stage.
  • We show that our approach outperforms state-of-the-art optimization based texture synthesis solutions on the Objaverse dataset. Score Distillation Sampling in pixel space allows omitting additional regularizers required in previous works. Additionally, we report the scaling performance of our method with the diffusion model size, the impact of lighting conditions used during training, and the role of physically based textures.
2 Related Work
Generating high-quality textures for 3D objects is a long-standing challenge in computer graphics and machine learning. Early approaches are based on category-specific generative models, such as generative adversarial networks (GANs) like Texture Fields [28], SPSG [11], LTG [42], and Texturify [36]. These methods often struggle with limited generalization and required extensive manual adjustments. Recent advances in large-scale text-to-image diffusion models [32, 33] transform texture synthesis, allowing more flexible and realistic results. However, different approaches vary in how they leverage diffusion models to incorporate 3D information and ensure global texture consistency.
Diffusion-based texture generation. A significant part of recent works formulates texture generation as an iterative inpainting process using depth-conditioned diffusion models. Text2Tex [6], TEXTure [31], and Paint3D [43] generate textures by synthesizing single-view images and progressively refining the texture using different object views. Although these zero-shot approaches allow for training-free generation, they often struggle with seams at visibility boundaries and suffer from baked-in lightning artifacts, leading to global inconsistencies. To mitigate these issues, some methods introduce additional refinement steps: Text2Tex employs an inpainting diffusion model for post-processing, while Paint3D refines textures directly in UV space. TexDreamer [22] follows a different approach by directly generating a UV texture map using a fine-tuned diffusion model, avoiding the need for iterative view synthesis and improving consistency across the surface.
An alternative strategy to address inconsistency is to introduce additional structures to enforce coherence between views. SyncMVD [21] tackles this by generating each view independently while employing a synchronization module that aligns outputs in every denoising step of the diffusion model. Similarly, TexFusion [5] suggest a module that aggregates multi-view object predictions into a unified latent texture representation, iteratively refining it throughout the diffusion process to ensure consistency across the entire object.
Another approach, inspired by MVDream [34], leverages consistent multi-view image generation as an initialization step for texture synthesis. Meta 3D TextureGen [2] follows this idea by generating four consistent views, which are projected onto a 3D surface to form an initial texture for subsequent refinement in UV space. MVPaint [8] extends this pipeline by introducing a spatial-aware 3D inpainting stage between projection and UV refinement, further enhancing coherence across different object views.
Optimization-based methods. Optimization-based techniques refine a texture by optimizing an objective function that evaluates rendered views. Early approaches such as CLIP-Mesh [26] and Text2Mesh [24] use differentiable rendering and CLIP [30] as a guidance signal to iteratively update texture parameters. However, these methods often produce blurry textures due to the lack of high-frequency details.
More recent works incorporate diffusion models to improve texture optimization. Latent-Paint [23], Fantasia3D [7], and Paint-it [41] combine differentiable rendering with the Score Distillation Sampling (SDS) technique [29], distilling knowledge from a diffusion model to refine textures. While this approach enhances realism, it often suffers from color oversaturation and high-frequency artifacts. FlashTex[13] mitigates some of these issues by employing a fine-tuned multi-view model to generate an initial textured view of the object, followed by an optimization stage using L2-loss and SDS for refinement. In this work, we continue the line of work employing SDS. However, we attribute some of the artifacts to the underlying latent diffusion model and propose using a pixel-space model. Aiming at physically-plausible textures, we use a text-to-image model pretrained on natural images and resort from using synthetic data for multi-view model adaptation.
PBR texture generation. Most of the works discussed above generate only RGB, yielding limited photo-realism. The works of Fantasia3D, Paint-it, and FlashTex propose to synthesize more advanced physically based textures [4], which are able to mimic different materials and look realistic under different lighting. We asses a similar approach to generation of PBR textures with a different generative model.
3 Preliminaries
In our work, we rely on a pretrained text-to-image model to generate textures. To apply a model designed for image generation to a different task, we use Score Distillation Sampling (SDS). Next, we describe the algorithm and the rendering algorithm required to apply it.
3.1 Score Distillation Sampling for Gradient Based Generation
SDS gained popularity in text-to-3D setup. In [29], the authors consider a differentiable rendering mechanism that produces a random view of a 3D model. In particular, they parameterize the model with a volumetric radiance field and render it using random lighting and camera position. SDS leverages a pretrained text-to-image diffusion model to guide the model parameters with gradient descent. Crucially, the mild assumptions on the rendering algorithm allow one to apply the method to other differentiable image generation mechanisms. In later works, Score-Distillation Sampling was successfully applied to other rendering functions and image parameterizations such as Gaussian Splatting [40, 9, 37], vector graphics [39, 14], and image rasterization with physically-based shading [20, 7, 41].
SDS defines gradients used to update the parameter $\theta\in\Theta$. Assume a mapping $x_{0}=g(\theta,z)$ that takes the parameter and a random variable $c$ and produces an image $x_{0}$. Given a diffusion process with noisy images $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ at time $t$ and scale-parameter $\bar{\alpha}_{t}$, the approach leverages a pretrained diffusion model $\epsilon_{\phi}(x_{t};y,t)$ to define a pseudo-objective used to update $x_{0}$. [^1] Throughout the paper, we assume that the diffusion model $\epsilon_{\phi}$ incorporates the condition $y$ through classifier-free guidance. The pseudo-objective $\hat{x}_{0}^{t}=\tfrac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\phi}(x_{t},t,y)}{\sqrt{\bar{\alpha}_{t}}}$ estimates the appearance of the rendered image $x_{0}$ based on a noisy sample $x_{t}$ with a single step diffusion model prediction.
The resulting gradient is estimated as follows:
$\nabla_{\theta}\mathcal{L}_{\text{SDS}}(\theta)=\mathbb{E}_{t,\epsilon,c}\left[w(t)(x_{0}-\hat{x}_{0}^{t})\frac{\partial x_{0}(\theta,c)}{\partial\theta}\right],$ (1)
where $w(t)$ is an adjustable scalar weight parameter. Note that [29] formulate the method in noise prediction $\epsilon$-terms, while we introduce the equivalent formulation in $x_{0}$-terms. On the one hand, the gradient is essentially the mean-squared loss gradient $\mathbb{E}\left[w(t)\|x_{0}-\hat{x}_{0}^{t}\|^{2}\right]$ with the omitted dependence of $\hat{x}_{0}^{t}$ on $\theta$. On the other hand, the analysis in [29] demonstrates that the procedure minimizes $\mathcal{L}_{SDS}(\theta)=\mathbb{E}_{t,c}\operatorname{KL}\left(q(x_{t}\mid t)\|p_{\phi}(x_{t}\mid y,t)\right)$ using the score $\epsilon_{\phi}(x_{t};y,t)=\nabla_{x_{t}}\log p_{\phi}(x_{t}\mid y,t)$.
Originally, SDS was applied to a pixel-space diffusion model that generates low-resolution images. However, pretrained pixel-space models are less common, and, as a result, applications of the method were primarily focused on latent diffusion models (LDMs). These models approximate the distribution of the low-dimensional embeddings of a high-resolution images rather than the images itself. Crucially, the mapping $\text{enc}(\cdot)$ that produces the embedding is differentiable, as a result, can be considered as a component of a function generating an image:
$g_{\textit{latent}}(\theta,c)=\text{enc}(g_{\textit{pixel}}(\theta,c)).$ (2)
The potential advantage of latent diffusion models is the ability to guide high-resolution images compared to pixel-space models. At the same time, it requires additional resources to run forward and backward passes of the encoder network. In Section 4, we analyze other implications of applying SDS in latent space rather than pixel space.
3.2 Differentiable Physically Based Shading
Our texture synthesis approach applies SDS in a setting, where $g(\theta,z)$ renders an image of a fixed 3D model given a set of texture parameters $\theta$ and a random camera position and scene lighting specified by $z$. For rendering, we follow the pipeline proposed by [27].
In particular, we take a differentiable rasterization algorithm based on the Nvdiffrast library by [19]. For textures, we adopt the Disney-physically based material model from [4], allowing us to incorporate the obtained textures into popular graphics engines without further modifications. The texture $\theta=(\theta_{d},\theta_{r},\theta_{m},\theta_{n})$ consists of a diffuse component $\theta_{d}\in\mathbb{R}^{h\times w\times 3}$, a roughness map $\theta_{r}$ parameterizing a specular lobe and a metalness factor $\theta_{m}$ used to interpolate between plastic and metallic appearance. Finally, a normal map $\theta_{n}$ stores surface normal information, allowing to simulate additional geometric detail even within low-polygon models.
To model lighting we incorporate cube environment map with a split sum approximation for the outgoing radiance integral approximation [16].
Figure 1: Overall pipeline. Our method consists of two stages. On the first stage, given a 3D mesh and a prompt our method employs a differentiable rendering pipeline to generate random views of the model under various lighting. We update the randomly initialized model texture using Score Distillation Sampling (SDS). On the second stage, refine the texture from the first stage using SDS with a super-resolution diffusion model. For the same view and lighting, we render two frames: a frame with the fixed texture taken from the first step and a frame with a current texture. Using the former frame as the condition, we back-propagate the SDS gradients through the latter frame.
Figure 2: Optimization based image generation for a toy single-mode diffusion model. We study the effect of image parameterization and the difference between pixel and latent space diffusion models. Score Distillation Sampling (SDS) finds the distribution mode for both latent and pixel space diffusion models. Top row: the pixel space model achieves near-perfect reconstruction quality measured by PSNR. The latent code $\text{enc}(\theta_{\text{RGB}})$ is also close to the mode and leads to high quality reconstruction. However, for the latent model the optimized parameter $\theta_{\text{RGB}}$ fails to approximate the original image. Bottom row: The implicit regularization proposed in [41] significantly improves the results for SDS for the latent diffusion model, but still falls short of the image reconstruction given by the decoder or pixel space diffusion model.
4 Methodology
In the following section, we start with a motivating example and then describe the proposed method. The overall scheme is illustrated in Figure 1.
4.1 Limitations of Score Distillation Sampling for Latent Diffusion Models
In this section, for latent diffusion models, we argue that the procedure is prone to producing samples with artifacts due to the ill-posed nature of the task.
Following the common practice, we convey the analysis on a task of image generation. Specifically, we use a mapping $g_{id}(\theta,c)=\theta$ that outputs a 2D image $\theta$ and does not rely on the second stochastic variable. The goal is to generate images on par with the diffusion model samples with a gradient descent algorithm. Intuitively, the SDS gradients guide images $g(\theta,c)$ towards the manifold of natural images. Using the chain rule, these gradients also propagate to the parameter $\theta$. However, in the case of latent diffusion models guidance happens within the latent space: $g(\theta,c)=\text{enc}(g_{id}(\theta,c))$. As a result, the training signal must propagate through the encoder.
Next, to mitigate some of the potential issues of SDS in our analysis, we consider a toy diffusion model that represents a degenerate distribution $\delta(x-x^{*})$ concentrated at $x^{*}$. The problems may include a poorly trained score estimation model or the excessive variance of the stochastic gradient 1. For the degenerate distribution, we use the optimal model that separates the Gaussian noise from the original sample
$\epsilon_{degenerate}(x_{t})=\tfrac{x_{t}-\sqrt{\bar{\alpha}_{t}}x^{*}}{\sqrt{1-\bar{\alpha}_{t}}}.$ (3)
Note that for the model from Eq. 3 and constant weights $w(t)=1$ the gradient in Eq. 1 simplifies to the deterministic gradient of the mean squared error:
$\nabla_{\theta}\mathcal{L}_{SDS}(\theta)=\nabla_{\theta}\|g(\theta)-x^{*}\|^{2}.$ (4)
In other words, in the specified toy setup, we can assume
$\mathcal{L}_{SDS}(\theta)=\|g(\theta)-x^{*}\|^{2}$ (5)
To analyze SDS, we ran the optimization in various setups and report the results in Figure 2. For the pixel-space diffusion model with $g(\theta)=\theta$, gradient descent rapidly converges to $\theta\approx x^{*}$ (top row, left). For the latent-space model we assume $x^{*}$ to be the latent code of the same image and minimize $\|\text{enc}(g(\theta))-x^{*}\|^{2}$. Although gradient descent converges and the decoder maps $\text{enc}(\theta)$ to the original image (top row, right), the optimized parameter $\theta$ itself is a poor approximation to the original image (top row, middle). In practice, SDS with latent diffusion models finds globally optimal solutions that fall outside of the natural image domain.
Notably, the above observation does not limit text-to-3D applications. However, in the above toy setup and texture synthesis domain the observed artifacts limit the applicability of SDS. To avoid the artifacts, a recent work [41] proposed to parameterize textures with a deep convolutional network. Their solution builds upon the observations of [38] and acts as an implicit regularizer. It helps avoiding solutions outside of the natural image domain, but requires running an additional convolutional and, as we show below, still leads to noticeable image artifacts.
In the bottom row of Figure 2, we report SDS performance with an implicit image parameterization from [41]. We denote the parameterization as DIP (Deep Image Prior), following [38]. In this case, the image is an output of a UNet network with weights $\theta_{dip}$ and a fixed input sampled from the standard Gaussian distribution. In the context of the SDS, we can interpret the forward pass of the UNet as the differentiable rendering algorithm $g_{\text{UNet}}(\theta_{dip})$ constrained to generate natural images.
We minimize Eq. 5 for the implicit parameterization assuming both latent and pixel-space models. The implicit parameterization significantly improves the result for the latent diffusion model (middle row). The artifacts present in the explicit parameterization are less noticeable. Still, even with the implicit parameterization SDS in the latent space is unable to match the reconstruction quality of the reconstructed latent code (middle and right bottom row images). In our preliminary experiments, we tried to impose additional regularization on $\theta$ using the decoder outputs, but did not achieve noticeable improvements compared to the default SDS. Besides that, pixel-space optimization achieves significantly better result for both explicit and implicit image parameterization.
Given the above observations, in our work, we use SDS with pixel-space models.
4.2 Applying SDS with Super-Resolution Diffusion Models
The downside of pixel-space diffusion models is that the model for high-resolution images require significant resources. While latent diffusion models employ a VAE [18] to encode a high-resolution image into a low-dimensional latent code, the typical solution for pixel-space models is to generate an image in a cascaded fashion. At the first stage, diffusion model generates a low-resolution image. During the second stage, a super-resolution model generates a high-resolution version of the image from the first stage.
In [29], the authors run SDS only with a low-resolution model. However, the low-resolution model may be unable to recover fine-grained details and high-frequency features as its training domain was limited to down-scaled images. To achieve highly detailed textures, in our work we mimic the architecture of cascaded diffusion models. First, we run SDS with a low-resolution diffusion model. Then, we refine the output of the first stage with SDS combined with a super-resolution model.
Figure 3: Score distillation sampling applied to a super-resolution diffusion model. Horizontal axis represents progress over time. Top row: the same image $g(\theta)$ used both as a diffusion model input $x$ and the condition $x_{cond}$ for super-resolution. The resulting procedure is unstable and significantly diverges from the original image. Bottom row: same procedure with fixed super-resolution condition $x_{cond}=g(\theta_{0})$. The condition anchors the dynamics to the initial image.
In text-to-image setup, the second stage model takes a low-resolution image as an input along with a text prompt and time. This leads to the two natural ways of applying SDS to refine the first stage output. The first option is to pass the rendered image $g(\theta,c)$ as the diffusion model argument in SDS and a down-scaled copy of the image as a condition. Intuitively, in this case, the algorithm gradually updates the texture by adding the details lacking after the previous gradient step. The second option is to freeze the parameter $\theta_{0}=\theta$ obtained at the first stage and run SDS using a copy of $\theta$ as the new parameter and the previous $\theta_{0}$ as a condition. In this case, the first stage output in the condition acts as an anchor for the gradient descent dynamics.
In our experience, the procedure in the first solution tends to be less stable and leads toward out-of-domain images, whereas the second approach tends to improve the resolution while preserving the details of the first stage result. We illustrate our observations in Figure 3.
In summary, we propose a texture synthesis approach based on a cascaded modification of the SDS. Figure 1 is a schematic representation of the overall approach.
5 Experiments
We evaluated our method on the Objaverse dataset [12], comparing it both qualitatively and quantitatively with state-of-the-art texture synthesis approaches. We also performed an ablation study to validate key design decisions.
5.1 Evaluation setup
For our method, we took DeepFloyd-IF [35] as an open-source cascaded diffusion model. We considered the XL ($64{\times}64$) and L ($256{\times}256$) models for the first and second stages, respectively. Also, we considered the setting with the smaller model IF(M) at the first stage to explore the effect of the diffusion model scale on the texture quality.
We compared the proposed method with three baselines that represent various approaches to texture synthesis. First, Text2Tex [6] is a well-established back-projection-based method generating non-PBR textures. Second, Paint-it [41] is a simplistic optimization-based method using SDS. The crucial distinction between our method and Paint-it is the underlying diffusion model and texture parameterization. Finally, FlashTex [13] is another recent approach using SDS. It addresses the limitations of previous methods by fine-tuning the model used for guidance on multi-view 3D data with controlled lighting. Again, all of the baselines are based on latent diffusion models for texture synthesis.
To evaluate the quality of generated textures, we computed the Frechet Inception Distance (FID) [15] and Kernel Inception Distance (KID) [3] metrics on a subset of the Objaverse dataset [12]. Specifically, we adopted the same curated subset used in Text2Tex [6], which includes 410 high-quality textured meshes from 225 distinct categories. This subset was originally filtered to exclude meshes with simplistic or inconsistent textures, category mismatches, and over-triangulated or scanned objects. All methods generated textures for the same set of untextured 3D meshes using identical text prompts, then each textured mesh is rendered from 20 fixed viewpoints at a resolution of $512\times 512$ with a white background using Blender [10]. We provide a detailed evaluation setup in Appendix A, where we specify the most critical rendering parameters (e.g., lighting configuration, render engine, etc.) to ensure accurate comparisons.
5.2 Main Results
Figure 4: Qualitative comparison. We synthesized textures for models from the Objaverse dataset using the proposed approach and a number of recent competing methods. Our method generates seamless textures with softer colors compared with latent diffusion-based approaches.
Figure 5: Human preference study comparing our method with competing optimization based and back-projection baselines.
Size FID KID Runtime
($\downarrow$) ($\downarrow$) (×$\times 10^{-3}$, $\downarrow$) (seconds, $\downarrow$)
Text2Tex 2.5B 21.3 4.1 549
Paint-it 1B 25.2 6.7 892
FlashTex 1.3B 29.0 9.5 280
CasTex (M) 400M 24.9 5.6 45
CasTex (M+L) 1.6B 23.2 5.0 628
CasTex (XL) 4.3B 22.8 4.4 174
CasTex (XL+L) 5.5B 20.5 3.1 710
Table 1: Comparison of FID and KID scores for different text-to-texture generation methods. Notably, even with the smallest diffusion model, our method outperforms optimization-based baselines. As expected, the super-resolution stage improves the results for both stages. Surprisingly, in our evaluation, the older back-projection-based method outperforms the optimization-based method and is only rivaled by our largest setup. We use NVIDIA A100 80GB for time measurements and generate textures with a resolution of 1024x1024 pixels.
Automated quality assessment. We report FID and KID scores in Table 1. Our smallest single-stage setup outperformed the two optimization-based approaches. The two-stage cascaded diffusion approach achieved the best performance, outperforming all baselines. We illustrate the two stages in Figure 6 and provide additional illustrations in Appendix C. The second stage significantly refines textures, enhancing fine-grained details and coherence.
Notably, the weak performance of FlashTex on the benchmark is the result of the limited generalization abilities of the Light ControlNet proposed by the authors. To our surprise, Text2Tex delivered competitive results. Even though the method is the oldest baseline, the back-projected photo-realistic images lead to photo-realistic textures and renders.
Our method also achieves competitive synthesis times, with the two-stage variant completing generation in 12 minutes per mesh – comparable to baselines while delivering better quality. For time-constraint applications, the single-stage variant generates plausible textures in just 3 minutes, offering a compelling trade-off between computational efficiency and texture realism, making it suitable for time-constraint applications.
User preference study. In addition, we conduct a human study to compare the perceptual quality of our generated textures against baseline methods. The evaluation is conducted by professional assessors, asked to make a decision between two rendered textures given a textual prompt. The decision is made according to the texture quality and textual alignment. For each pair, three responses are collected, and the final result is determined by majority voting. For evaluation, we use $256$ random objects from the previously mentioned Objaverse subset.
As shown in Figure 5, our two-stage approach outperforms all the baselines. Figure 4 presents a qualitative comparison of our method with the three baseline approaches. The comparison against our single-stage approach confirms that the second stage is highly effective for further texture refinement.
Also, we provide the results of our method using the smaller version of DeepFloyd-IF that contains 400M parameters. We observe that our method still outperforms competing works.
Overall, our method achieves impressive quality, effectively producing detailed, seamless, and high-fidelity textures, as further confirmed by quantitative results.
5.3 Ablation study
We conducted a series of experiments to understand the role of the key design choices in our approach.
FID KID
($\downarrow$) (×$\times 10^{-3}$, $\downarrow$)
CasTex (diffuse map only) 21.8 4.0
CasTex 20.5 3.1
Table 2: Ablation of Physically Based Rendering (PBR) components. Optimizing the full set of PBR components improves texture quality compared to synthesizing only the diffuse map.
Figure 6: Textures after the first and the second stage.
Do the PBR components improve the texture synthesis? PBR components enable the modeling of complex visual effects like reflections and material properties, avoiding baked-in lighting and improving realism.
We compare our full PBR-based approach with a baseline that generates only the diffuse map, one of the PBR components. As shown in Table 2, generating the full set of PBR maps leads to better FID and KID scores compared to generating the diffuse map alone. We provide visualization of PBR maps in Appendix E. Despite better metrics, a closer study of illustrations indicates that the method occasionally stuggles with baked-in lighting.
FID KID
($\downarrow$) (×$\times 10^{-3}$, $\downarrow$)
Text2Tex 21.3 4.1
CasTex (Text2Tex init, $k_{d}$ frozen) 21.3 4.2
CasTex (Text2Tex init, $k_{d}$ trained) 20.5 3.9
Table 3: Ablation study of initializing our second stage with Text2Tex outputs to generate PBR textures. The results show that allowing the diffuse map ($k_{d}$) to be optimized during our second stage leads to significant improvements in both FID and KID metrics, while keeping the diffuse map frozen maintains the same FID but slightly increases KID. This suggests that our approach can effectively serve as an enhancement module for existing texture generation methods.
Can our second stage improve other methods? We hypothesize that our second stage in the PBR texture generation pipeline can improve the results of other texture synthesis methods. To validate this, we experiment with applying our second stage to well-initialized diffuse maps from another method.
As shown in Table 3, our second stage can effectively generate PBR textures when initialized with diffuse maps from Text2Tex. We explore two configurations: (1) where we keep the initial diffuse map fixed ($k_{d}$ frozen) and (2) where we fine-tune the diffuse map during PBR generation ($k_{d}$ trained).
The results show that when the diffuse map is kept frozen, the FID score remains unchanged with a slight increase in KID. This may be because freezing $k_{d}$ also preserves any baked lighting information present in the initial diffuse map, preventing our method from correcting for these lighting inconsistencies. In contrast, when we allow diffuse map optimization ($k_{d}$ trained), our approach achieves significant improvements in both metrics, yielding the best FID and KID. This demonstrates that enabling diffuse map refinement allows our method to correct lighting artifacts and better harmonize the entire PBR material set, highlighting the potential of our approach as a versatile enhancement module for various texture generation pipelines.
FID KID
($\downarrow$) (×$\times 10^{-3}$, $\downarrow$)
Paint-it (Paint-it envlight) 25.2 6.7
Paint-it (proposed envlight) 24.5 6.4
CasTex (Paint-it envlight) 23.3 5.5
CasTex (proposed envlight) 20.5 3.1
Table 4: Ablation of environment lighting. We discovered that environmental lighting may play a crucial role in the plausibility of resulting physically-based textures. The default environmental lighting map in Paint-it presents an even gray lighting across all directions. The difference between a specular and a rough diffuse surface with such lighting is barely distinguishable. Instead, we propose an environmental lighting map with a single pronounced light source and augment lighting with horizontal rotations.
How does environment lighting affect results? Among recent works, Paint-it is the most similar to ours, as it also relies solely on SDS-based optimization. Unlike our approach, their rendering algorithm uses a brighter, uniform lighting environment map, whereas our setup combines a pronounced light source and a weaker ambient lighting component.
To understand the effect of environment lighting on SDS-based optimization, we evaluate both Paint-it’s lighting setup and our alternative configuration. As shown in Table 5, Paint-it achieves better results when using our lighting setup, slightly improving its FID and KID. Similarly, our methods performance degrades with Paint-it’s environment lighting. We hypothesize that, unlike uniform lighting, a single pronounce light source allows to highlight specular components of the PBR texture during training. The hypothesis and the above observations agree with the results from [13].
FID KID Runtime
($\downarrow$) (×$\times 10^{-3}$, $\downarrow$) (seconds, $\downarrow$)
CasTex (DIP, XL) 23.2 4.6 229
CasTex (XL) 22.8 4.4 174
CasTex (DIP, XL+L) 21 3.8 862
CasTex (XL+L) 20.5 3.1 710
Table 5: Deep Image Prior ablation. Even though DIP plays a crucial role in texture synthesis with latent diffusion models, for the cascaded diffusion models the induced regularization is obsolete and slightly reduces the resulting scores. Also DIP-based variants are more time-consuming, while all experiments generate textures at 1024$\times$1024 resolution.
DIP versus explicit parameterization. Following the approach of our closest competitor, Paint-it, we integrate Deep Image Prior (DIP) [38] into our method and evaluate its impact on texture optimization (see Section 4.1 for details). As shown in Table 5, our DIP-based variant achieves results comparable to our standard explicit parameterization approach but performs slightly worse in both FID and KID metrics while being significantly slower. Since explicit parameterization already delivers strong results and works faster, the added complexity of DIP does not offer a clear advantage in our framework.
6 Conclusion
In this work, we proposed an optimization-based texture synthesis approach based on cascaded diffusion models. We motivated the design with an analysis indicating the limitations of latent diffusion models. Our method improves upon recent optimization-based texture synthesis approaches in terms of automated metrics and user study. Even though our model benefits from using physically-based textures, we still observed baked-in lighting and occasional inconsistencies between a texture and the corresponding model geometry. Future work may address the baked-in lighting artifacts of the aforementioned limitations of Score Distillation Sampling with latent models.
Acknowledgements
We would like to thank Roman Chulkevich for his assistance with the computing cluster and Ksenia Ushatskaya for her help with the visualization of the generated textures used in the paper teaser. Special thanks go to Maxim Kodryan, whose mere existence made this research journey considerably more enjoyable. This research was supported in part through computational resources of HPC facilities at HSE University.
Appendix A Evaluation Protocol Details
A.1 Blender parameters
We render images using Blender 3.3.21 with the Cycles rendering engine. Eval camera positions are taken from a sphere with a radius of $1.4$, where the spherical coordinates $\varphi$ (azimuth angle) and $\theta$ (polar angle) are derived from the setup described in Text2Tex. The HDRI light map from polyhaven used in our experiments is visualized in Figure 7(a); we set the environment map strength to $0.7$.
Figure 7: Used HDRI light maps: (a) for evaluation; (b) for our pipeline; (c) for Paint-it.
A.2 Details of User Study
Every assessor was asked to evaluate the setup shown in Figure 8.
Figure 8: Assessor’s evaluation setup. Professional assessors were tasked with evaluating the quality of the generated textures and selecting the texture that best aligned with the suggested prompt. To ensure unbiased evaluations, the order of the methods was randomly shuffled for each setup.
Appendix B Environment lights
For our texture generation setup, we utilize the HDRI light map visualized in Figure 7(b). In contrast, the Paint-it HDRI light map, shown in Figure 7(c), represents an alternative lighting configuration. All the HDRI light maps used in this work are publicly available for download from polyhaven.
Appendix C Texture Synthesis Results in Verying Setups
In this section, we provide textures obtained with different diffusion model combinations to illustrate the impact of model size and the role of super-resolution model.
Figure 9: Textures obtained with middle (M) and extra-large (XL) diffusion models along with their versions improved with large (L) super-resolution model on the second stage.
We also provide additional illustrations to compare our method to the baselines.
Figure 10: Additional samples from the Objaverse dataset.
Appendix D High frequency study in DIP
Our investigation into integrating Deep Image Prior (DIP) with our framework reveals certain limitations that affect its practical utility in texture synthesis. Despite being computationally more expensive, our experiments suggest that DIP may not provide sufficient benefits to justify its implementation costs.
As shown in Figure 11, spectral analysis conducted on the Objaverse dataset [12] demonstrates why the parameterization of DIP performs poorly compared to our standard explicit approach. The spectral plot (Figure 11(a)) reveals DIP’s significant deficiency in high-frequency components, visually translating to outputs with noticeably less textural detail and surface variation. To further illustrate this effect, we provide additional examples in Figure 11(b) using two cow prompts, which clearly demonstrate the degradation of the detail in the generation based on DIP.
This limitation in creating detailed textures further validates our choice of explicit parameterization, which achieves better visual quality and quantitative metrics with less computational cost.
Figure 11: Spectral and visual comparison of texture generation approaches. (a) Power spectrum analysis (log scale) computed on the Objaverse subset, showing CasTex + DIP’s reduced power in high-frequency components compared to standard CasTex . (b) Additional examples showing texture generation results for two test prompts. CasTex produces rich, detailed textures while CasTex + DIP generates simplified outputs lacking fine-grained details, directly corresponding to its high-frequency limitations observed in the spectral analysis.
Appendix E Visualization of PBR Texture Maps
Even though physically based texture improve the overall results, our method occasionally struggles to fully disentangle the lighting effects. We visualize a few examples in Figure 12. For instance, consider partially baked highlights on the coffee maker or the suitcase.
Figure 12: Visualization of the PBR textures. Lighter colors indicate higher metallic and roughness values for the corresponding texutre components.